# FaithAct: Faithfulness Planning and Acting in MLLMs

### v1

@@ 5.1 Planning Objective
5.2 Planning Loop
At each iteration, FaithAct executes a faithfulness-first planning loop consisting of two sequential phases (as illustrated in Algorithm 1): (1) Poll and Ground for verification, and (2) Select and Abstain for action. The planner, a lightweight instruction-tuned MLLM, serves as a reasoning controller that composes executable steps such as Select, Abstain, or Count, conditioned on both textual inputs and the verified visual states. Each proposed step is immediately validated by the perceptual faithfulness metrics before execution, ensuring that the reasoning process evolves only through evidence-supported actions.
FaithAct provides a general and extensible interface composed of callable functions that supply multimodal evidence signals to guide the reasoning process. These functions act as structured APIs that the MLLM can invoke during planning, allowing the model to retrieve, verify, and reason over perceptual information in a unified manner:
  • Poll(): Returns the direct probability of an object’s existence based on the polling model (Section 4.2.1).
  • Ground(): Returns the bounding boxes and confidence scores of a claimed object detected by GroundingDINO (Section 4.2.2), providing spatial evidence.
  • Select(): Selects an object as existent if its overall confidence exceeds the threshold defined in Section 4.3.
@@ 5.2 Planning Loop
5.3 Refine-Based Reasoning
The overall reasoning process of FaithAct is summarized in Algorithm 1, which follows a refine-based procedure. After verification, any reasoning step that fails to meet the perceptual faithfulness threshold is not discarded outright but revised through a refine-based reasoning process. In this phase, the planner re-invokes the MLLM with updated evidence from the verifier, such as object existence labels, bounding boxes, and counts, to regenerate or adjust the unfaithful step. The refinement prompt explicitly conditions on verified visual information, guiding the model to produce a reasoning trace that remains consistent with perceptual evidence while maintaining logical continuity. This iterative correction mechanism ensures that FaithAct preserves both interpretability and grounding, progressively improving the reasoning chain through verified feedback. Further implementation and prompt details are provided in Appendix C.
6 Experiments
The purpose of this section is to illustrate the better performance of our planner on faithfulness of the reasoning chain, as well as good performance on correctness. We also include further experiments on ablation, distribution of $F_{step,{t}}$, etc. All results are reported in percentage (%).
Baseline Methods & Datasets RealWorldQA (%) MMHal (%) Average (%)
Prompt-based reasoning (no tools)
[…]
Tool-augmented reasoning (plan–act–observe)
[…]
Faithfulness-first planning (ours)
[…]
Table 1: Faithfulness evaluation results. We report mean and standard deviation of $F_{\text{chain}}$ in percentage (%). Bold and underline denote the best and second-best performance, respectively, within each model group.
Methods & Datasets RealWorldQA (%) MMHal (%)
Qwen-2.5-VL-7B+CoT 70.1 75.8
+FaithAct 74.5 76.8
Table 2: Correctness when applying FaithAct on Qwen-2.5-VL-7B. We choose the same benchmarks as Table 1.
Datasets. We select two real-world multimodal benchmarks RealWorldQA xai_grok1_5 and MMHal-Bench 2023llavarlhf (MMHal). The images in these benchmarks usually contain lots of real-world objects, and it’s convenient for faithfulness evaluation on them.
Models. We choose several widely-used MLLMs for our task. Qwen-2.5-VL-7B Bai2025Qwen25VLTR, InternVL3-8B Zhu2025InternVL3EA and LLaVA-one-Vision-1.5-8B (LLaVA-1.5-8B) an2025llavaonevision15fullyopenframework are chosen.
Baselines. We compare our method with several reasoning methods that do not training the MLLM itself. CoT wei2022chain; zhang2023multimodal is utilized as a baseline for simply generating reasoning chains, We also apply ReAct yao2022react, Grounded-CoT Wu2025GroundedCF, and Visual Abstract Thinking (VAT) liu2025visual for reasoning with the help of extra object information. ReAct enables MLLMs to automatically call useful functions during reasoning (we provide the same functions for it as our FaithAct). While Grounded-CoT requires MLLMs to ground the objects in reasoning by itself during reasoning, and VAT provide MLLMs abstract images (e.g., sketch) together with the original image to benefit from multi-level information.
Generation Hyper-parameters. The box threshold of GroundingDINO is 0.35 and the text threshold is 0.25. We keep the default generation settings for helper LLM and MLLMs.
6.1 Main Results and Findings
Faithfulness remains substantially underestimated. Across all evaluated models, faithfulness remains far from ideal. For instance, MLLMs such as Qwen-2.5-VL-7B achieve faithfulness scores of only around $\sim$50 across datasets and evaluation metrics, substantially below the desired level (near 100). This highlights that current reasoning models often generate partially ungrounded or inconsistent reasoning traces, indicating that perceptual verification and evidential grounding are still open challenges.
FaithAct improves faithfulness across models. Introducing the FaithAct framework consistently enhances reasoning faithfulness across three tested models. For example, InternVL3-8B achieves 57.35 $\pm$ 29.40 on RealWorldQA with FaithAct, compared to 44.23 $\pm$ 25.43 with CoT and 47.94 $\pm$ 19.36 with Grounded-CoT, respectively. Generally, our method attains the highest faithfulness in all the evaluated settings, demonstrating its effectiveness across different architectures and datasets. Here we provide a theoretical analysis on comparing FaithAct with ReAct in Appendix D and prove that FaithAct can perform better in faithfulness theoretically. Similar situations can be observed in the experiment results.
Figure 3: Distribution of average $F_{\text{step}}$ difference across reasoning steps. The x-axis are reasoning steps and y-axis represents the $F_{\text{step}}$ averaged difference between Qwen-2.5-VL-7B with and without FaithAct.
FaithAct mitigates hallucination. Our results show a significant improvement on hallucination-focused benchmarks, particularly on MMHal, where models integrated with FaithAct exhibit a marked reduction in hallucinated reasoning steps. Qwen-2.5-VL-7B with FaithAct achieves 66.45 $\pm$ 27.87, surpassing its second-best score of 56.77 $\pm$ 25.86. This notable gain confirms that evidence-first reasoning, as enforced by FaithAct, effectively constrains perceptual unfaithfulness during intermediate reasoning steps, thereby enhancing both local and global faithfulness of model outputs.
6.2 Faithfulness Without Performance Trade-off
Enforcing faithfulness does not degrade task performance. Table 2 reports the accuracy of Qwen-2.5-VL-7B with and without FaithAct across the three benchmarks [^2] We evaluate the scores ourselves due to the absence of official results in the technical report, serving as a baseline for future work.. The results show that integrating FaithAct preserves the model’s ability to generate correct final answers to multimodal questions, with slight improvements observed in two out of three datasets. These findings indicate that FaithAct enhances the faithfulness of reasoning steps without harming the model’s original performance. Moreover, although behavioral faithfulness is not explicitly measured, the empirical evidence supports our hypothesis that perceptually grounded reasoning encourages behavioral consistency.
Figure 4: Qualitative comparison of reasoning chains generated with and without FaithAct on two illustrative cases. In both tasks, FaithAct enforces step-level perceptual verification, correcting hallucinated descriptions (top in red) and producing more structured, visually grounded reasoning (bottom in blue).
6.2.1 Case Study
We conduct a qualitative comparison to illustrate how FaithAct improves reasoning faithfulness in the cases shown in Fig. 1. In the first example (Fig. 4 top), the baseline model hallucinates a yellow bicycle and no cars by relying on language priors or implication (e.g., yellow bus) rather than image evidence. With FaithAct, each reasoning step is perceptually verified, leading to the correct identification of a black bicycle and two cars. In the second example (bottom), both models correctly predict the next box in a visual sequence, but with FaithAct guidance, reasoning chain is more structured and explicitly justifies each visual attribute transition. For example, the direction of upward arrow in (C) is corrected to rightward. These cases demonstrate that FaithAct enhances perceptual grounding and visually consistent reasoning processes, empirically improving behavioral consistency.
6.3 Faithfulness Score Distribution
Fig. 3 reports the distribution of $F_{step,t}$, when comparing FaithAct and raw CoT without FaithAct. We observe that the benefit of FaithAct becomes particularly pronounced in the later reasoning steps, suggesting that its intervention is most effective when the model engages in deeper chains of reasoning. This observation is consistent with prior work wu2025more which points out that excessively CoT increases susceptibility to noise in their later steps and thereby leads to more hallucinations.
7 Conclusion
In this work, we investigated the problem of unfaithful reasoning in MLLMs. We first formalized two complementary notions, behavioral faithfulness and perceptual faithfulness, to disentangle reasoning consistency from evidential grounding. Building on this distinction, we introduced FaithEval, a systematic evaluation framework that quantifies faithfulness at both the step and chain levels. To move from evaluation to intervention, we proposed FaithAct that explicitly evaluates and requires evidential grounding at every reasoning step via lightweight callable functions before generating conclusions. Empirical results across multiple MLLMs demonstrate that models equipped with FaithAct outperform both prompt-based (CoT, VAT) and tool-augmented (Grounded-CoT, ReAct) baselines.
Despite these advances, several limitations still remain. FaithEval focuses primarily on object-level perceptual grounding, and may not fully capture higher-order relational or commonsense faithfulness. Moreover, FaithAct currently operates at the inference level; extending its principles into training-time optimization or fine-tuning objectives could further improve grounding efficiency. Finally, large-scale human or cognitive alignment studies are needed to validate whether perceptual faithfulness correlates with human trust and interpretability. We hope this work establishes faithfulness not merely as an evaluation objective, but as a core design principle for future multimodal reasoning systems that reason both correctly and evidentially.
Appendix A Object Extraction Prompts and Examples
[…]
Example 1:
Extract Result: [“coastal area”, “beach”, “city”]
Example 2:
Extract Result: [“taxis”, “buildings”]
Appendix B Training Details on POPE
Appendix C Prompts for FaithAct
[…]
Appendix D Proof of faithfulness by FaithAct v.s. ReAct
@@ Appendix C Prompts for FaithAct
$F_{\text{step}}\!\left(s_{g}^{\text{FaithAct}}\right)\geq F_{\text{step}}\!\left(s_{g}^{(0)}\right)=F_{\text{step}}\!\left(s_{g}^{\text{ReAct}}\right),\quad\forall g.$ (17)

### v3

@@ 4.1 Planning Objective
4.2 Planning-and-Acting Loop
FaithAct executes a faithfulness-first planning loop at each iteration. A lightweight instruction-tuned MLLM serves as the reasoning controller. This planner composes executable steps such as SELECT, ABSTAIN, or COUNT, based on both textual inputs and verified visual states. Each proposed step is immediately validated by perceptual faithfulness metrics before execution, ensuring that the reasoning process evolves only evidence-supported actions. To facilitate this, FaithAct provides an extensible interface composed of callable functions that serve as structured APIs. These functions supply multimodal evidence signals and can be invoked by the MLLM during planning, enabling the model to retrieve, verify, and reason over perceptual information in a unified and controlled manner.
Poll(): Returns the probability of objects’ existence based on the polling model (Sec. 3.2.1).
Ground(): Returns the bounding boxes and confidence scores of a claimed object detected by GroundingDINO (Sec. 3.2.2).
Select(): Selects an object as existent if its confidence score exceeds the threshold in Sec. 3.3.
@@ 4.2 Planning-and-Acting Loop
Baseline Methods & Datasets LLaVA-bench (%) RealWorldQA (%) POPE (%) MMHal (%) Average (%)
[…]
Table 1: Faithfulness evaluation across reasoning paradigms. We report the mean and standard deviation of the chain-level faithfulness score $F_{\text{chain}}$ (in %) on four benchmarks. Methods are organized by reasoning paradigm: prompt-based reasoning (gray shading), tool-augmented reasoning (blue shading), and faithfulness-first planning (green shading), while holding backbone models fixed to enable controlled comparison. Best and second-best results within each backbone model are highlighted in bold and underline, respectively.
4.3 Action-Guided Reasoning Refinement
The overall reasoning process of FaithAct is summarized in Algorithm 1 in Appx. E, which follows a refine-based procedure. After verification, any reasoning step that fails to meet the perceptual faithfulness threshold is not discarded outright but revised through an action-guided refinement process. In this phase, the planner re-invokes the MLLM with updated evidence from the evaluator, such as object existence labels, bounding boxes, and counts, to regenerate or adjust the unfaithful steps. The refinement prompt guides the model to produce a reasoning trace that remains consistent with perceptual evidence while maintaining logical continuity. This action-guided correction mechanism ensures that FaithAct preserves both interpretability and grounding, progressively improving the reasoning chain through verified feedback. Further implementation and prompt details are provided in Appx. D.
5 Experiments
We conduct an extensive experimental study demonstrating the effectiveness of Faithful-First RPA framework in measuring and improving faithfulness in MLLM reasoning. Unless otherwise noted, all results are reported as percentages (%).
Figure 3: Qualitative comparison of reasoning chains generated with and without FaithAct on two illustrative cases. In both tasks, FaithAct enforces step-level perceptual verification, correcting hallucinated descriptions (red colored in Fig. 1) and producing more structured, visually grounded reasoning (colored in blue).
Figure 4: Distribution of average $F_{\text{step}}$ difference across reasoning steps. The x-axis are reasoning steps and y-axis represents the $F_{\text{step}}$ averaged difference between Qwen with and without FaithAct.
5.1 Experimental Setup
Datasets. We evaluate our framework on widely used multimodal benchmarks covering object recognition, visual grounding, and hallucination-sensitive question answering, including LLaVA-bench Liu et al. (2023a), RealWorldQA xAI (2024), POPE Li et al. (2023b), and MMHal-Bench Sun et al. (2023) (MMHal). These datasets feature images with rich real-world objects and are well suited for evaluating perceptual faithfulness. Specifically, the queried objects in POPE are removed to ensure fair assessment.
Baselines. We compare our framework with representative training-free reasoning frameworks, including CoT Wei et al. (2022); Zhang et al. (2023), ReAct Yao et al. (2022), Grounded-CoT Wu et al. (2025a), and Visual Abstract Thinking (VAT) Liu et al. (2025). These methods are evaluated on widely used MLLMs, namely Qwen-2.5-VL-7B (Qwen) Bai et al. (2025), InternVL3-8B (InternVL) Zhu et al. (2025), and LLaVA-one-Vision-1.5-8B (LLaVA) An et al. (2025b).
Hyper-parameters setup. The box threshold of GroundingDINO is 0.35 and the text threshold is 0.25. We keep the default generation settings for helper LLM and MLLMs.
5.2 Experimental Results and Analysis
Perceptual faithfulness is broadly underestimated.
Across all evaluated models, faithfulness remains far from ideal. For instance, MLLMs such as Qwen achieve faithfulness scores of only around $\sim$50 across datasets and evaluation metrics, substantially below the desired level (near 100). This highlights that current reasoning models often generate partially ungrounded or inconsistent reasoning traces, indicating that perceptual verification and evidential grounding are still open challenges.
FaithAct improves faithfulness across models.
Introducing the FaithAct framework consistently enhances reasoning faithfulness across three tested models. For example, InternVL achieves 57.35 $\pm$ 29.40 on RealWorldQA with FaithAct, compared to 44.23 $\pm$ 25.43 without it. Generally, our method attains the highest faithfulness in 11 out of 12 evaluated settings, demonstrating its effectiveness across different architectures and datasets. Averaged across models, FaithAct achieves a mean score of 55.86%, outperforming the strongest baseline reasoning paradigm (ReAct, 48.10%) by 7.76 percentage points and CoT by 9.04%. We show that $F_{\text{chain}}$ of ReAct is theoretically bounded by that of FaithAct (in Appx. F). We also note that the reported standard deviations are relatively large, which is expected given the instance-level nature of perceptual faithfulness.
FaithAct mitigates hallucination.
Although FaithAct is not explicitly designed to mitigate hallucinations, it yields substantial gains on hallucination-focused benchmarks. In particular, on MMHal, models with FaithAct exhibit a marked reduction in hallucinated reasoning steps, achieving an average improvement of 21.99% and 9.81% over tool-augmented prompt-based reasoning frameworks, respectively. For example, Qwen with FaithAct achieves 66.45 $\pm$ 27.87, surpassing its second-best score. This notable gain indicates that the principle enforced by FaithAct effectively constrains perceptual unfaithfulness during intermediate reasoning steps, thereby enhancing faithfulness of model outputs.
Improving faithfulness does not degrade task performance.
Table 2 reports the performance of three models with and without FaithAct across the two benchmarks. The results show that integrating FaithAct preserves the model’s ability to generate correct final answers to multimodal questions, with slight improvements observed in two out of three datasets. These findings indicate that FaithAct enhances the faithfulness of reasoning steps without harming the model’s original performance. Moreover, although behavioral faithfulness is not explicitly measured, the empirical evidence supports our hypothesis that perceptually grounded reasoning encourages behavioral consistency.
Faithfulness increases across reasoning steps.
Fig. 4 reports the distribution of $F_{step}$, when comparing FaithAct and raw CoT without FaithAct. We observe that the benefit of FaithAct becomes particularly pronounced in the later reasoning steps, suggesting that its intervention is most effective when the model engages in deeper chains of reasoning. This observation is consistent with prior work Wu et al. (2025b) which points out that excessively CoT increases susceptibility to noise in their later steps and thereby leads to more unfaithfulness.
5.2.1 Qualitative Analysis and Case Studies.
We conduct a qualitative comparison to illustrate how FaithAct improves reasoning faithfulness in the motivating cases shown in Fig.1. In the first example (Fig. 3 top), the baseline model hallucinates a yellow bicycle and no cars by relying on language priors or implication (e.g., yellow bus) rather than image evidence. With FaithAct, each reasoning step is perceptually verified, leading to the correct identification of a black bicycle and two cars. In the second example (bottom), both models correctly predict the next box in a visual sequence, but with FaithAct guidance, reasoning chain is more structured and explicitly justifies each visual attribute transition. For example, the direction of upward arrow in (C) is corrected to rightward. These cases demonstrate that FaithAct enhances perceptual grounding and visually consistent reasoning processes, empirically supporting and improving behavioral consistency.
5.3 Human Validation on Extracted Objects
To evaluate whether Qwen extracted objects (Sec. 3.1) accurately capture what is indeed stated in the text, we conducted a human validation study using 50 snippet and 25 annotators, producing a total of $7,550$ object-level labels, detailed in Appx. ( G). Compared with human judgments, the LLM shows near-perfect consistency, achieving a precision of 99.42% with only $44$ false positives (over-extraction rate = 0.58%). At snippet level, we compute snippet validity, defined as the probability that all extracted objects for a snippet are judged correct by a human annotator. Among $1,250$ annotator–snippet pairs, the LLM attains a mean snippet validity of 0.9680, showing extracted object sets are almost fully aligned with human.
Datasets & Models RealWorldQA (%) MMHal (rating)
Qwen + CoT 70.1 3.40
+FaithAct (ours) 74.5 3.48
InternVL + CoT 70.8 3.61
+FaithAct (ours) 71.2 3.58
LLaVA + CoT 68.1 3.41
+FaithAct (ours) 67.8 3.46
Table 2: Task performance with CoT and FaithAct.
5.4 Ablation and Sensitivity Analysis
We conduct ablation and sensitivity analyses to examine the contribution of individual components in the proposed framework. In particular, we focus on the two core functions Poll() and Ground(). Throughout these experiments, the faithfulness evaluation protocol remains unchanged and we disable one module at a time to isolate its effect.
Results on Qwen are reported in Table 3. Removing either Poll() or Ground() leads to a noticeable decrease in faithfulness (approximately 5%), with Ground() having a slightly larger impact. This suggests that object localization provides essential visual evidence for perceptually grounded reasoning. The results indicate that the two modules are complementary, and that FaithAct achieves the greatest improvement in faithfulness when both are jointly enabled.
We additionally conduct a study replacing GroundingDINO with a recently published alternative, SAM3 Carion et al. (2025). Equiped with SAM3, the performance of FaithAct drops obviously (5% on RealWorldQA and more on MMHal). This may suggest that FaithAct needs localization-specific models (like GroundingDINO) to provide grounding results more accurate with proper confidence.
Figure 5: Comparative performance of the FaithAct framework using SAM3 vs. GroundingDINO as the Ground() function. Results show mean accuracy ($\%$) and standard deviation (error bars) across three MLLMs on the RealWorldQA and MMHal datasets.
Datasets & Models RealWorldQA (%) MMHal (%)
[…]
Table 3: Ablation study of the two core components.
6 Conclusion
We introduced Faithful-first RPA, a framework that enforces perceptual grounding throughout the reasoning process. Within this framework, FaithEvi provides a principled and fine-grained evaluation of perceptual faithfulness, while FaithAct operationalizes faithfulness-first reasoning through planning and acting. Experiments across benchmarks show that our framework improves perceptual faithfulness by up to 24% without compromising accuracy, effectively mitigating hallucinations. These results underscore the value of faithfulness as a core design principle. Future work will extend to behavioral faithfulness and more challenging open-ended reasoning settings.
Limitations
This work primarily focuses on perceptual faithfulness and does not directly evaluate behavioral faithfulness, i.e., the alignment between reasoning traces and the model’s final decision process. While our empirical results suggest that enforcing perceptual grounding may be associated with more behaviorally consistent outputs, this relationship is not explicitly measured or guaranteed.
In addition, although we conduct a human validation study indicating that LLM-extracted objects are generally accurate, we do not perform large-scale human evaluations of step-level or chain-level perceptual faithfulness. Complementary human studies at different levels of granularity could help further contextualize and validate our findings.
Finally, our current implementation verifies perceptual faithfulness primarily at the level of object existence. Extending the Faithful-first RPA framework to incorporate attribute- and relation-level verification remains an important direction for future work, and may further reduce perceptual unfaithfulness in cases where objects are present but their properties or relations are mischaracterized.
Appendix A Related Work
Reasoning Frameworks for MLLMs. MLLMs have developed surprising reasoning capabilities in multimodal problem solving, task planning, scientific discovery and so on Lu et al. (2024); Li et al. (2025a); Gao et al. (2024); An et al. (2024, 2025a); Lin et al. (2025); Li et al. (2025b, c). To enhance systematic reasoning, several frameworks Chen et al. (2024a); Sun et al. (2025); Shao et al. (2024); Zhou et al. (2025); Wang et al. (2025) have been proposed to decompose multimodal problems into interpretable steps. Among them, CoT Chen et al. (2024a), Grounded-CoT Wu et al. (2025a) and ReAct frameworks Yao et al. (2022) are famous and insightful ones. More recently, VAT Liu et al. (2025) explores hierarchical and compositional reasoning, encouraging models to form abstract visual concepts that support complex decision-making. Such generated reasoning chains, though often fluent and logically structured, may still include steps unsupported by visual evidence or inconsistent with the model’s actual decision process Yu et al. (2024).
Faithfulness in Multimodal Reasoning. Despite the demonstrated effectiveness of reasoning in improving the performance of MLLMs, an unfaithful problem emerges: the reasoning traces generated by these models do not always behave faithfully to the internal processes that produce their final answers Lyu et al. (2023); Turpin et al. (2023); Lanham et al. (2023). Models often rely on latent knowledge or shortcut associations that are not explicitly expressed in their reasoning chains. As a result, the generated reasoning steps may read as a plausible but untrustworthy explanation, motivating the need for explicit faithfulness assessment Jacovi and Goldberg (2020); Parcalabescu and Frank (2023); Matton et al. (2025); Barez et al. (2025); Arcuschin et al. (2025). Several works have introduced benchmarks or metrics to evaluate faithfulness explicitly. M4 Li et al. (2023a), FaithScore Jing et al. (2023) and TIFA Hu et al. (2023) propose metrics for evaluating the faithfulness for vision-language models. However, we find that behavioral alignment (behavioral faithfully) does not guarantee the correctness of the final output (see left panel in Fig.1).
Object Hallucination as Unfaithful Consequences. Object hallucination is identified as a common challenge, as a consequence, in Large Vision-Language Models (LVLMs), where the model describes or reasons about objects absent in the input image. Several benchmarks, such as POPE Li et al. (2023b) and MMHal-Bench Sun et al. (2024), have been developed to systematically evaluate this phenomenon. However, these efforts typically treat hallucination as an isolated failure mode rather than as a manifestation of broader unfaithful reasoning.
Recent studies address hallucination through different mechanisms: (i) training-time alignment, such as hallucination-aware Reinforcement Learning from Human Feedback or preference optimization Sun et al. (2024); Zhang (2024); (ii) decoding-time constraints, including grounded or contrastive decoding Li et al. (2023b); and (iii) feature-level grounding strategies that enhance cross-modal alignment Ghosh et al. (2025). Together, these advances underscore that faithful reasoning requires not only linguistic coherence but also perceptual accountability, further motivating to verify evidential grounding before inference.
Appendix B Object Extraction Prompts and Examples
[…]
Example 1:
Extract Result: [“coastal area”, “beach”, “city”]
Example 2:
Extract Result: [“taxis”, “buildings”]
Appendix C Training Details on POPE
Appendix D Prompts for FaithAct
[…]
Appendix E Algorithm of FaithAct
Here we list the algorithm process of FaithAct in Algorithm 1.
@@ Appendix E Algorithm of FaithAct
Appendix F Proof of faithfulness by FaithAct v.s. ReAct
@@
$F_{\text{step}}\!\left(s_{g}^{\text{FaithAct}}\right)\geq F_{\text{step}}\!\left(s_{g}^{(0)}\right)=F_{\text{step}}\!\left(s_{g}^{\text{ReAct}}\right),\forall g.$ (15)
@@
Appendix G Human Validation of LLM-Extracted Objects
This section provides detailed methodology and analysis for the human validation study used to assess the accuracy of LLM-extracted objects described in Sec. 3.1.
G.1 Study Design and Data Collection
We randomly sampled 50 text snippets from our evaluation corpus. Each snippet was processed by the Qwen to extract a set of candidate objects that were intended to represent entities explicitly or implicitly stated in the text. To validate the correctness of these extracted objects, we conducted a human annotation study involving 25 annotators.
Each annotator independently evaluated all extracted objects for each snippet. For every object, annotators answered the binary question: “whether the extracted object explicitly exists in the text.” Objects were labeled as supported (1) if they were explicitly mentioned or unambiguously implied by the text, and as unsupported (0) otherwise. Annotators were instructed to rely solely on the provided text and to avoid using external world knowledge. When uncertain, they were instructed to mark the object as unsupported.
In total, the study yielded 7,550 object-level labels, corresponding to 1,250 annotator–snippet pairs (50 snippets $\times$ 25 annotators).
G.2 Evaluation Metrics
We first evaluate precision at the object level by comparing LLM-extracted objects against human judgments. An extracted object is considered a false positive if it is labeled as unsupported by a human annotator. We compute object-level precision as
$\text{Precision}=\frac{\text{\# supported objects}}{\text{\# extracted objects}}.$
Across all object-level annotations, the LLM achieves a precision of 99.42%, with only 44 false positives, corresponding to an over-extraction rate of 0.58%.
While precision metrics capture local correctness, they do not reflect whether an entire set of extracted objects for a snippet is jointly accurate. To assess holistic correctness, we introduce snippet validity, defined as an indicator function over annotator–snippet pairs:
$\text{SV}(a,s)=\mathbf{1}\big[\forall o\in\mathcal{O}_{s},\;o\text{ is supported by annotator }a\big],$
where $\mathcal{O}_{s}$ denotes the set of objects extracted for snippet $s$.
We report mean snippet validity by averaging $\text{SV}(a,s)$ across all annotator–snippet pairs:
$\text{Mean SV}=\frac{1}{|\mathcal{A}||\mathcal{S}|}\sum_{a\in\mathcal{A}}\sum_{s\in\mathcal{S}}\text{SV}(a,s).$
Across 1,250 annotator–snippet pairs, the LLM achieves a mean snippet validity of 0.97. This indicates that for nearly all annotator–snippet evaluations, all extracted objects for a snippet are judged correct.
Summary
The combined object-level and snippet-level analyses provide complementary views of the choice of Qwen as an object extractor. High object-level precision and high snippet validity demonstrates that extracted object sets are almost always entirely accurate, serving as reliable inputs for downstream faithfulness evaluation in our framework.
Figure 6: Human annotation interface for object-existence validation. For each snippet, annotators are presented with the original text and a set of objects automatically extracted by the LLM. Annotators judge whether each object is explicitly mentioned or unambiguously implied by the text, producing binary Supported (1) or Unsupported (0) labels used for evaluating extraction faithfulness.